What is the solution for:
X4 + 11X2 - 80
and
2a2 - 11a + 12

Answer by checkley75(3666)

X^2+11X-80=0
(X-5)(X+16)=0
X-5=0
X=5 ANSWER.
X+16=0
X=-16 ANSWER.